y=4x-2
y=-7x+5
-2y=12x+18

Answer by Alan3354(69443)

y=4x-2
y=-7x+5
-2y=12x+18
------------
There are 3 eqns in 2 variables. Might not be a common solution.
---
y=4x-2
y=-7x+5
---
4x-2 = -7x + 5
11x = 7
x = 7/11
--
y = 6/11
-------
Check the 3rd eqn
-12/11 =? 84/11 + 18
-96/11 <> 18
No solution for all 3.
=================================
If you graph the 3 lines, you get 3 intersections.
There's no point that all 3 pass thru.